The area under the standard normal curve where p( 0 < z < 1.31)


1
Expert's answer
2021-07-16T08:18:11-0400
"P(0<Z<1.31)=P(Z<1.31)-P(Z\leq 0)"

"\approx0.904902-0.5=0.404402"